Mr.Reyes is 20 years older than his son, how old are the father and the son if Mr. Reyes is a 6 times as old as his son?
Answer by stanbon(75887)

Mr.Reyes is 20 years older than his son,
R = s + 20
-------------
how old are the father and the son if Mr. Reyes is a 6 times as old as his son?
R = 6s
-------------
Solve by substitution:
6s = s+20
5s = 20
s = 4 yrs old (Son's age)
R = s+20 = 4+20 = 24 yrs old (Mr. Reyes age)
================
Cheers,
Stan H.